OpenStudy (anonymous):

Approximate the solution to 1/sqrt{x^2+} = 1/x+5

OpenStudy (anonymous):

Possibilities: (a) x = -2.8320 (b) x = -1.2062 (c) x = -2.4000 (d) x = -1.2000 (e) x = 0.3846

OpenStudy (anonymous):

o it should be 1/sqrt{x^2+1}

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

is that \[\frac{1}{\sqrt{x^{2}+1}} = \frac{1}{x+5} \]

OpenStudy (anonymous):

if so just cross multiply (or invert) and square both sides then solve for x to get \[x = -12/5 =-2\frac{2}{5} = -2.4\]

OpenStudy (anonymous):

could you write out the steps please?

OpenStudy (anonymous):

I listed the steps

OpenStudy (anonymous):

ok so i get x^2+1=sqrt{x+5} then?

OpenStudy (anonymous):

Almost. You would get sqrt(x^2 + 1) = x + 5

OpenStudy (anonymous):

ok i got that but then u said square both sides so that would get rid of the square on one side and place it on the other

OpenStudy (anonymous):

Not exactly. When you square both sides that means to multiply it by itself or raise it to the power of 2 meaning \[(\sqrt{x^{2}+1})^{2} = (x+5)^{2} \]

OpenStudy (anonymous):

yes i know that but when u raise the side with the radical to the power of 2, they cancel out

OpenStudy (anonymous):

correct, but then the term on the other side is squared (not put under a radical as u had said)

OpenStudy (anonymous):

ok